OpenStudy (anonymous):

Which of the following statements best describes the graph of –5x + 2y = 1? It is a curve joining the points (–5, 2), (2, 3), and (4, 1). It is a curve joining the points (–1, –3), (–1, –3), and (1, 5). It is a straight line joining the points (1, 3), (3, 8), and (–3, –7). It is a straight line joining the points (4, –3), (–1, 2), and (–4, 5).

OpenStudy (anonymous):

Start by solving for Y \[y=1/2+(5/2)x\] Since x is a linear term(that is, its exponent is 1) we know that the graph is linear since we solved for y. Now it's a matter of eliminating C or D. The easiest way to do this is by "plugging in" the points into the equation. Let's try (1,3) y=1/2+(5/2)*3=1/2+15/2=16/2=8 Immediately, we know that it cannot be C because the point (1,3) doesn't lie on the line--it doesn't satisfy the equation. The only remaining answer is D. Does that make sense?

OpenStudy (anonymous):

perfect sense. Thank you so much(: